# Lost in the FoG: Pitfalls of Models for Large-Scale Hydrogen Distributions

### v4

@@ Appendix A Resolution Effects
Figure 12: Comparisons of the ${\rm H}$i (brown) and galaxy (blue and red) biases from the primary volumes of TNG100 (dark colors) and TNG300 (light). The galaxy biases agree between the two resolutions, particularly at early redshifts and for red galaxies. Conversely, the TNG100 ${\rm H}$i bias is vertically offset by $\sim$25from TNG300, although their scale-dependencies agree. Our predictions of model errors are robust against the vertical offset because they are not dependent on the precise value of the ${\rm H}$i bias.
We study the resolution dependence of our results in Fig. 12 by comparing the ${\rm H}$i, blue, and red galaxy biases from TNG100 and TNG300, which have resolutions that differ by approximately an order of magnitude (Nelson et al., 2019). We find that red galaxy bias is nearly identical between the two resolutions within overlapping scales, and the blue galaxy bias converges for $z=1$. At later redshifts, the blue galaxy bias has a slight vertical offset, similar to the ${\rm H}$i bias at all redshifts. The different behavior between the various tracers is expected (Section 6.1) as the blue or ${\rm H}$i-rich galaxies tend to have masses closer to the resolution limit than red galaxies, particularly at late redshifts (Villaescusa-Navarro et al., 2018; Nelson et al., 2018). Furthermore, star-formation rates are generally enhanced with better resolution, leading to less ${\rm H}$i overall (Davé et al., 2013).
Both the correlation coefficients and pairwise velocity dispersions (PVD) are also impacted significantly by resolution. The effects on the correlation coefficient can be inferred from Fig. 12 — the downward shift of the ${\rm H}$i bias implies a decrease in the ${\rm H}$i auto power spectrum, which causes all ${\rm H}$i-galaxy correlation coefficients to shift down as well (online figures). Similarly, all PVDs are expected to be dependent on resolution due to changing satellite fractions altering how virialized motions manifest in redshift space (Slosar et al., 2006). Given these resolution effects, we cannot conclude that the bias, correlation coefficient, and PVD values presented in Section 4 are converged with resolution.
However, our conclusions about model performance in Section 5 are independent of resolution effects. The precise values of the large-scale bias and correlation coefficient do not impact model error estimates, since they are removed in the $P_{\mathrm{model}}/P_{\mathrm{HI}}$ ratios. Tracer RSDs can change at different mass resolutions due to how they sample the large-scale velocity field (Jennings et al., 2015; Robertson et al., 2024), but we find similar RSD errors in TNG100 VN18.
@@ Appendix B Shot Noise
Shot noise arises in power spectra due to the discrete sampling of a continuous density field. Generally, this term is thought to be negligible on large scales but can be significant on small scales (Castorina & Villaescusa-Navarro, 2017). We establish the role of shot noise in our results by first outlining its expected behavior and then comparing bias definitions with and without shot noise in Fig. 13.
@@
Here, $V$ is the volume and $M_{i}$ is the mass of the $i$th source, which in our case would be an ${\rm H}$i mass element or a galaxy. If the shot noise is purely Poisson, then $P_{\mathrm{SN}}$ is scale-independent, although $P_{\mathrm{SN}}$ can become slightly scale-dependent via residual correlations on the smallest scales (Chen et al., 2021).
Instead of folding the shot noise term into the bias like in Equation 2, we can separate the two terms with an alternative bias definition:
@@
We note that our fiducial models (Equations 12-13) and models like Equation B3 fundamentally include the same components, but attribute the shot noise to different terms. Consequently, our tests of the models using scale-dependent biases remain the same regardless of bias definition, although our analysis of the errors associated with the bias term change. We compare our fiducial bias definition to the bias from Equation B2 in Fig. 13 and re-interpret the errors linked to the bias in the context of this alternative bias definition.
@@ Appendix C Redshift Evolution
All models in both the auto and cross power spectra perform worse at later redshifts due to stronger nonlinearities and baryonic effects. Both effects cause tracer biases to become more scale-dependent at larger scales and complicate how RSDs affect the clustering. This is particularly true at $z=0$, when assuming a constant ${\rm H}$i bias introduces significant errors into the corresponding models (cB, cB+FoG, tK+cB, and tK+cB+FoG). The RSD errors in sB+FoG and tK+sB+FoG clearly manifest uniquely at $z=0$ as compared to the other redshifts. Attributing this phenomenon to any one process is difficult, although we note that the ${\rm H}$i-halo mass relation and ${\rm H}$i density profiles change relatively rapidly between $z=1$ and $z=0$ (see figures 4 and 5 from Villaescusa-Navarro et al., 2018).

### v5

@@ Appendix A Resolution Effects
Figure 12: Comparisons of the $\rm H$}i (brown) and galaxy (blue and red) biases from the primary volumes of TNG100 (dark colors) and TNG300 (light). The galaxy biases agree between the two resolutions, particularly at early redshifts and for red galaxies. Conversely, the TNG100 $\rm H$}i bias is vertically offset by $\sim$25% from TNG300, although their scale-dependencies agree. Our predictions of model errors are robust against the vertical offset because they are not dependent on the precise value of the $\rm H$}i bias.
We study the resolution dependence of our results in Fig. 12 by comparing the $\rm H$}i, blue, and red galaxy biases from TNG100 and TNG300, which have resolutions that differ by approximately an order of magnitude (D. Nelson et al., 2019). We find that red galaxy bias is nearly identical between the two resolutions within overlapping scales, and the blue galaxy bias converges for $z=1$. At later redshifts, the blue galaxy bias has a slight vertical offset, similar to the $\rm H$}i bias at all redshifts. The different behavior between the various tracers is expected (Section 6.1) as the blue or $\rm H$}i-rifch galaxies tend to have masses closer to the resolution limit than red galaxies, particularly at late redshifts (VN18, D. Nelson et al., 2018). Furthermore, star-formation rates are generally enhanced with better resolution, leading to less $\rm H$}i overall (R. Davé et al., 2013).
Both the correlation coefficients and pairwise velocity dispersions (PVD) are also impacted significantly by resolution. The effects on the correlation coefficient can be inferred from Fig. 12 — the downward shift of the $\rm H$}i bias implies a decrease in the $\rm H$}i auto power spectrum, which causes all $\rm H$}i-galaxy correlation coefficients to shift down as well (online figures). Similarly, all PVDs are expected to be dependent on resolution due to changing satellite fractions altering how virialized motions manifest in redshift space (A. Slosar et al., 2006). Given these resolution effects, we cannot conclude that the bias, correlation coefficient, and PVD values presented in Section 4 are converged with resolution.
However, our conclusions about model performance in Section 5 are independent of resolution effects. The precise values of the large-scale bias and correlation coefficient do not impact model error estimates, since they are removed in the $P_{\mathrm{model}}/P_{\mathrm{HI}}$ ratios. Tracer RSDs can change at different mass resolutions due to how they sample the large-scale velocity field (E. Jennings et al., 2015; A. Robertson et al., 2024), but we find similar RSD errors in TNG100 ( VN18, ).
@@ Appendix B Shot Noise
Shot noise arises in power spectra due to the discrete sampling of a continuous density field. Generally, this term is thought to be negligible on large scales but can be significant on small scales (E. Castorina & F. Villaescusa-Navarro, 2017). We establish the role of shot noise in our results by first outlining its expected behavior and then comparing bias definitions with and without shot noise in Fig. 13.
@@
Here, $V$ is the volume and $M_{i}$ is the mass of the $i$th source, which in our case would be an $\rm H$}i mass element or a galaxy. If the shot noise is purely Poisson, then $P_{\mathrm{SN}}$ is scale-independent, although $P_{\mathrm{SN}}$ can become slightly scale-dependent via residual correlations on the smallest scales (Z. Chen et al., 2021).
Instead of folding the shot noise term into the bias like in Equation 2, we can separate the two terms with an alternative bias definition (see also equation 7 from R. Ando et al., 2019):
@@
We note that our fiducial models (Equations 12-13) and models like Equation B3 fundamentally include the same components, but attribute the shot noise to different terms. Consequently, our tests of the models using scale-dependent biases remain the same regardless of bias definition, although the errors we attribute to the bias term would change. We compare our fiducial bias definition to the bias from Equation B2 in Fig. 13 and re-interpret the errors linked to the bias in the context of this alternative bias definition.
@@ Appendix C Redshift Evolution
All models in both the auto and cross power spectra perform worse at later redshifts due to stronger nonlinearities and baryonic effects. Both effects cause tracer biases to become more scale-dependent at larger scales and complicate how RSDs affect the clustering. This is particularly true at $z=0$, when assuming a constant $\rm H$}i bias introduces significant errors into the corresponding models (cB, cB+FoG, tK+cB, and tK+cB+FoG). The RSD errors in sB+FoG and tK+sB+FoG clearly manifest uniquely at $z=0$ as compared to the other redshifts. Attributing this phenomenon to any one process is difficult, although we note that the $\rm H$}i-halo mass relation and $\rm H$}i density profiles change relatively rapidly between $z=1$ and $z=0$ (see figures 4 and 5 from VN18, ).